# thanks: ORCID: $0$-0003-2020-0803thanks: ORCID: $0$-0001-7285-0707thanks: ORCID: $0$-0003-0665-266X

### v1

@@ III.3 Testing model misspecification
In the previous section, we demonstrated the utility of the KL matrix as a diagnostic tool for tracking the evolution and stability of network training. In this section, we will further develop this idea and explore the possibility of using the KL matrix to detect model misspecification. Once again, the underlying hypothesis is simple: given the flexibility of neural network models, we have no guarantees on the extrapolation behaviour of a trained model beyond data in the training set. In particular, for an ensemble of models, we cannot expect that if network predictions agree (as quantified by low values in the KL matrix, for example) within the training sample that they will each extrapolate in an identical manner. This is broadly a statement about the large degeneracy that exists in the parameter space of large neural networks when optimising a scalar loss function. In general, this is presented as a negative feature, however, here we point out that one can actually leverage this fact to detect model misspecification.
@@
and standard deviation $\sigma_{k}=\sigma=0.01$. The parameter $\theta_{i}\in[-1,1]$ gives a constant contribution to the mean of the dataset and
@@
For the example presented in this section, we fix $N_{d}=10000$ as the size of our training dataset [^4] To be more accurate, as explained in appendix A, we draw $N_{d}$ values for $\theta$, which we keep fixed during the training process and, every time we show a pair $(\theta,d_{k})$ to the networks in the ensemble, we draw an independent realisation of $d_{k}$ given $\theta$. Following this procedure, we effectively show many instances of the data corresponding to the same parameters, which allows the networks to better learn the statistical properties of the data. and fix $a=0$ for all examples used during the training process. The ensemble used for this example consists of $N_{\rm Net}=5$ networks. After fixing the network architecture and training strategy, we train and monitor the evolution of the KL divergence among all networks in the ensemble. At the end of our run, the maximal value of the KL divergence is $\sim 7\times 10^{-2}$, corresponding to a maximal mismatch of .$\sim.37\sigma∼ .$ in the peak position (see section II.3 for the conversion from KL values to $\sigma$-levels). As already mentioned, for all datasets we used during the training process, we fix $a=0$, meaning we only train our networks on examples that do not contain the contribution due to the bump. In other words, if the observed data consisted solely of $\theta$, the candidate model would be correctly specified. This establishes a clear reference point against which the effects of misspecification can be assessed. Specifically, we can compare the observed values in the KL matrix for a given (potentially misspecified) observation and compare them to $7\times 10^{-2}$, or look for posteriors that differ by more than .$\sim.37\sigma∼ .$.
@@
In Fig. 4, we compare the results obtained conditioning the networks of the ensemble on an observation matching data used for the training process (left plot), i.e., $a=0$, and an observation that is misspecified (right plot) with $a=3\times 10^{-2}$. The two plots in the first row show the data realisation with and without adding the bump component. It is evident that with the addition of the bump, the data deviates from the simulation model (which assumes a flat mean across the bins) used during the training process. In the bottom row, we show the posteriors obtained from the different networks in the ensemble after conditioning on the corresponding observation. Specifically, on the left hand side, which excludes the bump, all posteriors agree well (i.e., to the level expected from the KL matrix analysis) with each other and with the true posterior (shown in black). On the other hand, on the right hand side, we see a much larger spread among the posteriors given by the different networks. To be more quantitative, after conditioning on an observation that lies outside the training set, we find a KL divergence among the networks of the ensemble that significantly exceeds the threshold value used during the training process. We highlight this visually by showing the orange vertical contours, which denote the ±$\pm 0.37\sigma$ variance expected from the in-sample training result. We can see that the posteriors in the misspecified model disagree to a level that is not captured by this error.
@@ IV Outlook and conclusions
Moreover, we demonstrated in Section III that ensemble disagreement, as measured by the KL divergence, can serve as a practical detector for model misspecification. Conditioning trained estimators on out-of-distribution observations results in a measurable increase in divergence between ensemble members, underscoring the diagnostic power of this approach beyond mere convergence assessment. Our experiments with controlled perturbations to observed data confirmed that the KL divergence can track systematic deviations beyond the information provided by the loss function. It would be interesting in future work to benchmark the sensitivity of this test of model misspecification against classical techniques or other ML-enhanced tools. It would also be interesting to explore further ways to enhance ensemble learning for SBI. For example, by drawing inspiration from parallel tempering, training models with different learning rates, allowing occasional information exchange via the KL matrix could improve robustness and generalisation, as well as achieving a better balance between exploration and exploitation. Furthermore, in analogy with Reversible Jump MCMC (RJ-MCMC), dynamically adapting model architectures within an ensemble would allow models to “jump” between different complexities based on performance, potentially leading to more efficient use of computational resources. Beyond leveraging ensembles for consistency checks, combining outputs from diverse models also offers a natural path to more stable posterior estimates by mitigating variance across architectures.
@@ Appendix A Further details and tests
Figure 6: Training and Validation Losses. Training (light colour) versus validation loss (full colour) for the different dataset sizes (100, 200, 500, 1000 top row, left to right, and 2000, 5000, 10000, 20000 left to right bottom row) as a function of epochs for all the networks in the ensemble (red, gree, blue) discussed in Section III.2.
@@
We conclude this appendix by providing further details on the models used in Section III.3. As already mentioned in the main text, for this example, parameters $\theta$ are uniformly sampled from $[-1,1]$. The corresponding observations $x$ are produced by a simulator that adds Gaussian noise ($\sigma=0.01$) to $\theta$ across the 10 independent measurements. In practice, each observation corresponds to a 10-dimensional data vector modelling noisy measurements. The ensemble used for this example consists of five networks trained in parallel using the procedure described above in this appendix. The training dataset consists of 10000 $\theta$ samples (the observations are resampled dynamically as discussed above), and the validation dataset has size 2000. In the left panel of Fig. 7, we show the training/validation losses for the five networks in the ensemble as a function of training epochs. As expected, after a steep initial decrease, the losses of the five networks flatten out and fluctuate randomly around an asymptotic value. In the right panel of the same figure, we show the evolution of the KL divergences among the different networks in the ensemble as functions of training epochs. Again, after an initial phase, they all tend to scatter around some nearly constant value. We notice that all KL divergences involving network ID 1 are significantly larger, indicating that there is less consistency between this particular model and the others in the ensemble. For reference, the maximum value of KL among all combinations at the end of the training process is $\simeq 7\times 10^{-2}$. We use this number in the main text as the lower bound to which we compare detections of model misspecification.

### v2

@@ III.3 Testing model misspecification
In the previous section, we demonstrated the utility of the KL matrix as a diagnostic tool for tracking the evolution and stability of network training. In this section, we will further develop this idea and explore the possibility of using the KL matrix to detect model misspecification. "Misspecification" in our framework refers to the scenario where the simulator used for training does not accurately represent the data-generating process for the real observed data. This mismatch is revealed only when we apply the trained estimators to actual observations, as during training the model is correctly specified by definition. The underlying hypothesis is simple: given the flexibility of neural network models, we have no guarantees on the extrapolation behaviour of a trained model beyond data in the training set. In particular, for an ensemble of models, we cannot expect that if network predictions agree (as quantified by low values in the KL matrix, for example) within the training sample that they will each extrapolate in an identical manner. This is broadly a statement about the large degeneracy that exists in the parameter space of large neural networks when optimising a scalar loss function. In general, this is presented as a negative feature, however, here we point out that one can actually leverage this fact to detect model misspecification.
A crucial prerequisite for this test is to distinguish between training-induced variance and misspecification-induced variance. We define two quantities: (i) the training-time KL divergence, $\mathrm{KL}_{\mathrm{train}}$, which is measured on simulated test data from the same distribution as the training data, and quantifies the baseline disagreement due to the training process. (ii) The inference-time KL divergence, $\mathrm{KL}_{\mathrm{obs}}$, measured on the actual observed data. Our misspecification diagnostic is based on the difference $\Delta=\mathrm{KL}_{\mathrm{obs}}-\mathrm{KL}_{\mathrm{train}}$. A significant positive value of $\Delta$ suggests the observation is out-of-distribution, as it triggers enhanced disagreement among ensemble members compared to their baseline. Therefore, the test is most powerful when the ensemble has achieved a low, stable baseline $\mathrm{KL}_{\mathrm{train}}$, otherwise any signal from misspecification could be lost in the noise of training variance.
@@
and standard deviation $\sigma_{k}=\sigma=0.01$. The parameter $\theta$ is drawn from a uniform prior $\theta\sim\text{Uniform}(-1,1)$ and gives a constant contribution to the mean of the dataset and
@@
For the example presented in this section, we fix $N_{d}=10000$ as the size of our training dataset [^3] To be more accurate, as explained in appendix A, we draw $N_{d}$ values for $\theta$, which we keep fixed during the training process and, every time we show a pair $(\theta,d_{k})$ to the networks in the ensemble, we draw an independent realisation of $d_{k}$ given $\theta$. Following this procedure, we effectively show many instances of the data corresponding to the same parameters, which allows the networks to better learn the statistical properties of the data. and fix $a=0$ for all examples used during the training process. The ensemble used for this example consists of $N_{\rm Net}=5$ networks. After fixing the network architecture and training strategy, we train and monitor the evolution of the KL divergence among all networks in the ensemble. At the end of our run, the maximal value of the KL divergence is $\sim 7\times 10^{-2}$, corresponding to a maximal mismatch of $\sim.37\sigma$ in the peak position (see section˜II.3 for the conversion from KL values to $\sigma$-levels). As already mentioned, for all datasets we used during the training process, we fix $a=0$, meaning we only train our networks on examples that do not contain the contribution due to the bump. In other words, if the observed data consisted solely of $\theta$, the candidate model would be correctly specified. This establishes a clear reference point against which the effects of misspecification can be assessed. Specifically, we can compare the observed values in the KL matrix for a given (potentially misspecified) observation and compare them to $7\times 10^{-2}$, or look for posteriors that differ by more than $\sim.37\sigma$.
@@
In Fig. 4, we compare the results obtained conditioning the networks of the ensemble on an observation matching data used for the training process (left plot), i.e., $a=0$, and an observation that is misspecified (right plot) with $a=3\times 10^{-2}$. The two plots in the first row show the data realisation with and without adding the bump component. It is evident that with the addition of the bump, the data deviates from the simulation model (which assumes a flat mean across the bins) used during the training process. In the bottom row, we show the posteriors obtained from the different networks in the ensemble after conditioning on the corresponding observation. Specifically, on the left hand side, which excludes the bump, all posteriors agree well (i.e., to the level expected from the KL matrix analysis) with each other and with the true posterior (shown in black). On the other hand, on the right hand side, we see a much larger spread among the posteriors given by the different networks. To be more quantitative, after conditioning on an observation that lies outside the training set, we find a KL divergence among the networks of the ensemble that significantly exceeds the threshold value used during the training process. We highlight this visually by showing the orange vertical contours, which denote the $\pm 0.37\sigma$ variance expected from the in-sample training result. We can see that the posteriors in the misspecified model disagree to a level that is not captured by this error.
@@
Classical approaches to model misspecification detection include goodness-of-fit tests based on $\chi^{2}$ statistics comparing observed and expected data distributions, posterior predictive checks, and Bayesian model comparison via evidence ratios. In our specific example with the bump model, a $\chi^{2}$ test on the residuals between the data and the best-fit flat model would indeed detect the misspecification at comparable sensitivity levels ($\approx 1-2\sigma$). While classical tests like the $\chi^{2}$ statistic are computationally cheaper and should be a standard first check, our ensemble-based approach offers several complementary advantages: (i) it does not require explicit specification of a null hypothesis or alternative model, (ii) it provides a continuous diagnostic that can be monitored during analysis rather than requiring a separate validation step, and (iii) it naturally accounts for parameter uncertainty through the ensemble disagreement rather than relying on point estimates. Recent ML-enhanced approaches include using classifier-based test statistics Cranmer et al. (2016), where a neural network is trained to distinguish between data generated under different hypotheses. Our KL-based approach is conceptually related but operates in posterior space rather than data space, making it naturally integrated with the inference procedure itself. A systematic comparison of these approaches across various misspecification scenarios would be valuable future work.
@@ IV Outlook and conclusions
Moreover, we demonstrated in Section III that ensemble disagreement, as measured by the KL divergence, can serve as a practical detector for model misspecification. Conditioning trained estimators on out-of-distribution observations results in a measurable increase in divergence between ensemble members, underscoring the diagnostic power of this approach beyond mere convergence assessment. Our experiments with controlled perturbations to observed data confirmed that the KL divergence can track systematic deviations beyond the information provided by the loss function. It would be interesting in future work to perform a systematic comparison of the sensitivity of this KL-based test against classical techniques and other ML-enhanced tools. Another interesting direction would be to deliberately train on misspecified simulations (e.g., with known simplifications) and use the KL diagnostic to assess when and where these simplifications become problematic. It would also be interesting to explore further ways to enhance ensemble learning for SBI. For example, by drawing inspiration from parallel tempering, training models with different learning rates, allowing occasional information exchange via the KL matrix could improve robustness and generalisation, as well as achieving a better balance between exploration and exploitation. Furthermore, in analogy with Reversible Jump MCMC (RJ-MCMC), dynamically adapting model architectures within an ensemble would allow models to “jump” between different complexities based on performance, potentially leading to more efficient use of computational resources. Beyond leveraging ensembles for consistency checks, combining outputs from diverse models also offers a natural path to more stable posterior estimates by mitigating variance across architectures.
@@
Data Availability
The code to reproduce all results in this manuscript is publicly available on GitHub at https://github.com/james-alvey-42/sbi-ensemble-diagnostics, and the datasets generated and analysed during the current study are available in the Zenodo repository, https://zenodo.org/records/17522623.
@@ Appendix A Further details and tests
The Gaussian Mixture Model example used in Section III.2 follows the default configuration from the sbibm benchmark library Lueckmann et al. (2021) to ensure reproducibility. The specific model details are as follows:
  • The prior distributions for both parameters are uniform: $\theta_{1},\theta_{2}\sim\text{Uniform}(-10,10)$.
  • The likelihood $p(x|\theta)$ is a mixture of two 2D Gaussian distributions with means at $(\theta_{1},\theta_{2})$ and $(-\theta_{1},-\theta_{2})$, with equal weights of 0.5.
  • The standard deviations of each mode are as specified in the sbibm benchmark library implementation Lueckmann et al. (2021).
Figure 6: Training and Validation Losses. Training (light colour) versus validation loss (full colour) for the different dataset sizes (100, 200, 500, 1000 top row, left to right, and 2000, 5000, 10000, 20000 left to right bottom row) as a function of epochs for all the networks in the ensemble (red, green, blue) discussed in Section III.2.
@@
Figure 8: Evolution of maximum/average KL divergences. Behaviour of the maximum, minimum and average values of the KL divergences in the ensembles trained for the example discussed in Section III.2. We see that all three quantities decrease systematically as the dataset size increases until convergence is reached.
@@
We conclude this appendix by providing further details on the models used in Section III.3. As already mentioned in the main text, for this example, parameters $\theta$ are uniformly sampled from [1,1][-1,1]. The corresponding observations $x$ are produced by a simulator that adds Gaussian noise ($\sigma=0.01$) to $\theta$ across the 10 independent measurements. In practice, each observation corresponds to a 10-dimensional data vector modelling noisy measurements. The ensemble used for this example consists of five networks trained in parallel using the procedure described above in this appendix. The training dataset consists of 10000 $\theta$ samples (the observations are resampled dynamically as discussed above), and the validation dataset has size 2000. In the left panel of Fig. 7, we show the training/validation losses for the five networks in the ensemble as a function of training epochs. As expected, after a steep initial decrease, the losses of the five networks flatten out and fluctuate randomly around an asymptotic value. In the right panel of the same figure, we show the evolution of the KL divergences among the different networks in the ensemble as functions of training epochs. Again, after an initial phase, they all tend to scatter around some nearly constant value. We notice that all KL divergences involving network ID 1 are significantly larger, indicating that there is less consistency between this particular model and the others in the ensemble. For reference, the maximum value of KL among all combinations at the end of the training process is $\simeq 7\times 10^{-2}$. We use this number in the main text as the lower bound to which we compare detections of model misspecification.